# M2IST: Multi-Modal Interactive Side-Tuning for Efficient Referring Expression Comprehension

### v2

@@ Abstract
Referring expression comprehension (REC) is a vision-language task to locate a target object in an image based on a language expression. Fully fine-tuning general-purpose pre-trained vision-language foundation models for REC yields impressive performance but becomes increasingly costly. Parameter-efficient transfer learning (PETL) methods have shown strong performance with fewer tunable parameters. However, directly applying PETL to REC faces two challenges: (1) insufficient multi-modal interaction between pre-trained vision-language foundation models, and (2) high GPU memory usage due to gradients passing through the heavy vision-language foundation models. To this end, we present M2IST: Multi-Modal Interactive Side-Tuning with M3ISAs: Mixture of Multi-Modal Interactive Side-Adapters. During fine-tuning, we keep the pre-trained uni-modal encoders fixed, updating M3ISAs on side networks to progressively connect them, enabling more comprehensive vision-language alignment and efficient tuning for REC. Empirical results reveal that M2IST achieves an optimal balance between performance and efficiency compared to most full fine-tuning and other PETL methods. With M2IST, standard transformer-based REC methods present competitive or even superior performance compared to full fine-tuning, while utilizing only 2.11% of the tunable parameters, 39.61% of the GPU memory, and 63.46% of the fine-tuning time required for full fine-tuning.
@@ I Introduction
Referring expression comprehension (REC) is one of the most challenging vision-language tasks, aiming to locate a specific object in an image based on a given referring expression [1, 2, 3, 4, 5, 6]. Recent studies [3, 7, 8, 9] have demonstrated impressive performance by fine-tuning general-purpose vision-language foundation models for this task. However, fully fine-tuning these pre-trained models is computationally expensive when adapting to a new dataset (see Figure 1 (a)). Additionally, fine-tuning on limited REC data may lead to catastrophic forgetting and overfitting, as recently evaluated in other vision-language tasks [10, 11].
@@
Recently, parameter-efficient transfer learning (PETL) methods [12, 13, 14, 15] have been proposed to address similar issues by updating only a small set of parameters to efficiently adapt pre-trained models to downstream tasks. Adapter-tuning [12], a prominent PETL method, has demonstrated significant success across various vision-language foundation models as well as downstream tasks [11, 16, 17]. It typically inserts a tunable lightweight bottleneck-shaped module sequentially into each frozen backbone layer. Most transformer-based REC models [3, 7, 18, 19] use pre-trained Vision Encoder and Language Encoder to separately extract image and text features, which are then integrated to form multi-modality features for reasoning. A straightforward approach to apply adapter-tuning for REC is to insert the adapters into the transformer encoder layers to enhance fine-tuning efficiency (see Figure 1 (b)). However, this introduces two significant challenges: (1) Updating inserted adapters still requires backpropagation of gradients through the large pre-trained vision and language models, placing a heavy burden on GPU memory during fine-tuning (see Figure 1 (b)). (2) Vision-language foundation models are pre-trained separately, each with its own structure and training data [20, 21]. Directly inserting vanilla adapters into them may lack cross-modality interaction in the shallow layers of the entire modal, resulting in sub-optimal vision-language alignment. This is particularly problematic for predicting referred objects in cases with complex semantic information, such as human actions and spatial relations (see Figure 4 "Without Interaction").
@@
To address these challenges, we propose a novel Multi-Modal Interactive Side-Tuning (M2IST) method that effectively strengthens vision-language alignment and enables parameter- and memory-efficient transfer to REC within the unified interactive side networks (see Figure 1 (c)). Specifically, we introduce Mixture of Multi-Modal Interactive Side-Adapters (M3ISAs), which incorporate Vision Expert Adapters (VEA), Language Expert Adapters (LEA), and Interaction Expert Adapters (IEA) into the side networks in parallel with the heavy encoders. VEA and LEA transfer pre-trained single-modality knowledge to the REC domain. IEA utilizes a linear layer for weight-sharing between image and text features, enabling progressive interaction between the referring sentence and input image. Such interaction aggregates channel-level vision-language alignment at shallow layers of the model, facilitating deep token-level vision-language alignment in deeper layers for improved performance. This elegant design achieves parameter-, memory-, and time-efficient intra- and inter-modality representation transfer for REC.
@@
We conduct extensive experiments on RefCOCO [22], RefCOCO+ [22], and RefCOCOg [23, 24] to demonstrate the effectiveness and efficiency of M2IST for REC. Experimental results show that M2IST achieves the optimal performance-parameter-memory trade-off compared to most full fine-tuning methods and other PETL methods. By applying our M2IST method, a standard transformer-based REC model can requiring only 2.11% of the tunable parameters, 39.61% of the GPU memory and 63.46% of the fine-tuning time compared to full fine-tuning, while still achieving competitive performance (see Figure1). With the sufficient vision-language interaction strengthened by our M3ISAs, our method can accurately locate the referred objects for various complex cases (see Figure 4).
@@ II-A Referring Expression Comprehension
Referring expression comprehension (REC) [1, 2, 3, 25, 7, 26, 4, 9] aims to locate the specific objects in images based on textual descriptions.
Early REC methods [1, 27, 28] follow a two-stage pipeline that first uses a pre-trained object detector (e.g., Faster R-CNN [29]) to generate a set of sparse object proposals, which are then ranked by their similarity to the given textual description. MAttNet [1] uses Faster R-CNN to generate object proposals, then scores them based on the referring expression to find the most relevant object. RvG-Tree [30] builds a relational visual graph to capture object interactions and ranks proposals according to their relevance to the expression, improving grounding accuracy. However, these two-stage REC methods heavily rely on the quality of the object proposals and cannot directly predict the referred object region. One-stage anchor-based methods [2, 31, 32, 33] have been developed to eliminate the proposal generation step, directly predicting object bounding boxes from predefined anchors [34]. FAOA [2] utilizes the YOLOv3 detector [34], integrating it with an encoded language vector to ground the referred regions. MRLN [4] introduces three modules: feature-feature relational learning, feature-task relational learning, and task-task relational learning, to enhance the collaborative learning of REC, effectively reducing prediction inconsistency in multi-task learning. Recently, transformer-based methods [3, 35, 26, 7, 9] have demonstrated superior performance by implicitly modeling cross-modality relationships within a unified architecture. The pioneering work TransVG [3] applies a stack of transformer encoders to perform feature extraction and multi-modal fusion for REC. VGTR [35] employs a transformer encoder-decoder architecture to jointly reason over visual and textual inputs, grounding the referred object without relying on pre-trained detectors or word embeddings. Dynamic M-DETR [9] uses a dynamic multi-modal transformer decoder to adaptively sample visual features and perform text-guided decoding for REC.
As REC models continue to scale up, their performance has improved. However, this performance gain comes at the cost of increased computational cost, which demands larger GPU memory for more parameter fitting (see Figure 1 (a)).
@@ II-B Parameter-efficient Transfer Learning
Recent PETL methods can be classified into two categories. The first category is updating additional parameters in modules inserted into the model [12, 15, 36] or appended to the input data [14, 37, 38]. Adapter [12] incorporates a bottleneck module into each Transformer layer, positioned after both the Multi-Head Attention (MHA) and the Feed-Forward Networks (FFN). AdaptFormer [15] embeds the adapter module parallel to the FFN in each encoder of a Vision Transformer. VPT [14] appends learnable visual vectors to the input sequences (VPT-Shallow) or to the input of each transformer encoder layer (VPT-Deep). The second category involves decomposing weight matrices into two low-rank matrices and updating only the small factorization matrices [13, 39]. As a pioneering work, LoRA [13] integrates a tunable pair of low-rank decomposed weight matrices into each encoder layer of the pre-trained networks. FacT [39] incorporates tunable factorized weight matrices into each layer of the pre-trained networks. There is also growing interest in adapter-based PETL methods for vision-language tasks, including text-image retrieval [11] and text-video retrieval [40, 41]. Most of these methods aim to achieve effective cross-modality interaction while maintaining parameter efficiency.
However, existing PETL methods still face substantial GPU memory consumption during the fine-tuning stage, as gradients must propagate through the heavy pre-trained encoders for REC (see Figure 1 (b)).
@@ II-C Memory-efficient Transfer Learning
In this work, inspired by the existing efforts, our M2IST bridges the pre-trained Vision Encoder and Language Encoder through the unified side interactive networks, enabling a parameter-, memory-, and time-efficient transfer to the REC task (see Figure 1 (c)).
@@ III-A Base Architecture
We apply a standard transformer-based REC model as our base architecture, shown in Figure 1 (a), which comprises: (1) a Vision Encoder, (2) a Language Encoder, and (3) a Vision-language Encoder.
@@ III-B Multi-Modal Interactive Side-Tuning: M2IST
Given that the pre-trained vision and language encoders contain rich knowledge and comprise about 95% of the model’s parameters. We first explore two approaches to reduce training overhead:
@@
To address the above issues, we propose Multi-Modal Interactive Side-Tuning (M2IST) that keeps the pre-trained encoders frozen and updates the proposed Mixture of Multi-Modal Interactive Side Adapters (M3ISA) on side networks to facilitate parameter- and memory-efficient fine-tuning for REC, as shown in Figure 2. Note that we do not show the LayerNorm for simplicity.
@@
The core component of our M2IST is M3ISA (see Figure 2 (right)), which consists of two distinct adapters, i.e., intra- and inter-modality adapters to effectively and efficiently bridge the pre-trained Vision Encoder and Language Encoder.
The intra-modality adapters include Vision Expert Adapter (VEA) and Language Expert Adapter (LEA), shown as separate blue branch and green branch in Figure 2 (right). They adhere to a fundamental design [12] for transferring pre-trained single-modality representations to more domain-specific ones. Specifically, both of them consist of a down-projection layer $\mathbf{W}_{\text{down}}$, ReLU non-linear activation, and an up-projection layer $\mathbf{W}_{\text{up}}$ in sequence. Taking the VEA as a formulaic example, given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, the function of VEA can be formally expressed as:
@@
In the base architecture, the Vision Encoder and Language Encoder are responsible for extracting single-modality features, while only the V-L encoder is tasked with cross-modality fusion through token-level vision-language alignment. However, such the late fusion has been proved insufficient when the referring sentence contains complex semantic information, such as spatial relationships [47]. Therefore, we introduce an inter-modality adapter, named the Interaction Expert Adapter (IEA). Specifically, each IEA shares a set of tokens within the same channel dimensions $C_{i}$ (i.e., Interacted Tokens in Figure 2 (right)) between the Vision Encoder and Language Encoder, thereby achieving channel-level vision-language alignment during the multi-modal feature extraction stage. As depicted by the entire pink section in Figure 2 (right), IEA include a unique down-projection layer for vision $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.180,0.459,0.714}C_{v}}\times C_{d}}$ and language $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.439,0.678,0.278}C_{l}}\times C_{d}}$, ReLU activation, an interactive up-projection layer $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times{\color[rgb]{0.929,0.420,0.380}C_{i}}}$, and a unique up-projection layer for vision $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.180,0.459,0.714}C_{v}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$ and language $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.439,0.678,0.278}C_{l}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$, where $C_{v}$, $C_{l}$, and $C_{i}$ represent the vision, language, and interaction channels, respectively. Given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and language tokens $\bm{x}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$, the corresponding down-projection layers first down-sample them to the bottleneck features $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{d}}$ and $\bm{z}_{l}\in\mathbb{R}^{N_{l}\times C_{d}}$. Then, the corresponding up-projection layers and interactive up-projection layer up-sample these bottleneck features and concatenate them within the same modality to obtain the cross-modality features $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$ as:
@@
Our IEA enables lightweight adaptation of cross-modality representations without increasing the explicit computational burden of vision-language feature extraction, thereby efficiently bridging the pre-trained vision and language encoders. Our experimental results in Table IX further demonstrate that deeper channel-level vision-language alignment yields better performance for REC. Moreover, such alignment will further improve the token-level vision-language alignment in V-L Encoder, leading to better region-sentence understanding in challenging REC scenarios, as analyzed in Section IV-E.
@@
As depicted in Figure 2 (left), we incorporate a stack of M3ISAs into two side networks that operate in parallel with the pre-trained dual encoders. Specifically, in one encoder layer (both for vision and language), the IEA first receives processed vision/language tokens from the Multi-head Attention (MHA) layers as input and produces adapted, interacted tokens for the vision/language side network. Subsequently, the VEA/LEA take the processed vision/language tokens from the Feed Forward Networks (FFN) as input and generate adapted single-modality tokens for the corresponding side networks. The outputs of the IEA and VEA/LEA are added within the vision/language side networks, along with the original vision/language tokens through skip-connections. After passing through the side networks, the outputs of the vision/language side networks are added to the outputs of the vision/language encoders. During fine-tuning, we keep the pre-trained encoders fixed and update the M3ISAs in the side networks, allowing the pre-trained encoders to act as standalone feature extractors.
@@
Following most transformer-based REC methods [3, 26, 9], the training loss function is a combination of the widely used smooth L1 loss and GIoU loss. Specifically, the prediction is donated as $\mathbf{b}=(x,y,w,h)$, and the normalized ground-truth box as $\hat{\mathbf{b}}=(\hat{x},\hat{y},\hat{w},\hat{h})$. The training objective is:

### v3

@@ Abstract
Referring expression comprehension (REC) is a vision-language task to locate a target object in an image based on a language expression. Fully fine-tuning general-purpose pre-trained vision-language foundation models for REC yields impressive performance but becomes increasingly costly. Parameter-efficient transfer learning (PETL) methods have shown strong performance with fewer tunable parameters. However, directly applying PETL to REC faces two challenges: (1) insufficient multi-modal interaction between pre-trained vision-language foundation models, and (2) high GPU memory usage due to gradients passing through the heavy vision-language foundation models. To this end, we present M2IST: Multi-Modal Interactive Side-Tuning with M3ISAs: Mixture of Multi-Modal Interactive Side-Adapters. During fine-tuning, we fix the pre-trained uni-modal encoders and update M3ISAs to enable efficient vision-language alignment for REC. Empirical results reveal that M2IST achieves better performance-efficiency trade-off than full fine-tuning and other PETL methods, requiring only 2.11% tunable parameters, 39.61% GPU memory, and 63.46% training time while maintaining competitive performance. Our code is released at https://github.com/xuyang-liu16/M2IST.
@@ I Introduction
Referring expression comprehension (REC) is one of the most challenging vision-language tasks, aiming to locate a specific object in an image based on a given referring expression [1, 2, 3, 4, 5, 6]. Recent studies [3, 7, 8, 9] have demonstrated impressive performance by fine-tuning general-purpose vision-language foundation models for this task. Following this trend but with increasingly larger model sizes, early approaches such as TransVG [3], Dynamic M-DETR [9] adapt pre-trained DETR [10] and BERT [11], while more recent works scale up to larger architectures - MaPPER [12] employs DINOv2 [13], and SimVG [14] and C3VG [15] utilize pre-trained BEiT-3 [16] to achieve better performance for REC. Although these approaches demonstrate continuous performance improvements, fully fine-tuning pre-trained models remains computationally expensive when adapting to new datasets (see Fig. 1 (a)). Additionally, fine-tuning on limited REC data may lead to catastrophic forgetting and overfitting, as recently evaluated in other vision-language tasks [17, 18, 19].
@@
Recently, parameter-efficient transfer learning (PETL) methods [20, 21, 22, 23] have been proposed to address similar issues by updating only a small set of parameters to efficiently adapt pre-trained models to downstream tasks. Adapter-tuning [20], a prominent PETL method, has demonstrated significant success across various vision-language foundation models as well as downstream tasks [18, 24]. It typically inserts a tunable lightweight bottleneck-shaped module sequentially into each frozen backbone layer. Most transformer-based REC models [3, 7, 25, 26, 27] use pre-trained Vision Encoder and Language Encoder to separately extract image and text features, which are then integrated to form multi-modality features for reasoning. A straightforward approach to apply adapter-tuning for REC is to insert the adapters into the transformer encoder layers to enhance fine-tuning efficiency (see Fig. 1 (b)). However, this introduces two significant challenges: (1) Updating inserted adapters still requires backpropagation of gradients through the large pre-trained vision and language models, placing a heavy burden on GPU memory during fine-tuning (see Fig. 1 (b)). (2) Vision-language foundation models are pre-trained separately, each with its own structure and training data [10, 11]. Directly inserting vanilla adapters into them may lack cross-modality interaction in the shallow layers of the entire modal, resulting in sub-optimal vision-language alignment. This is particularly problematic for predicting referred objects in cases with complex semantic information, such as human actions and spatial relations (see Fig. 4 "Without Interaction"). Most recently, a series of memory-efficient transfer learning (METL) methods [28, 29, 30, 31] have been proposed to reduce GPU memory consumption during fine-tuning. However, these METL methods are primarily designed for single-modality tasks, such as image recognition [28, 29], dense prediction [31], and natural language understanding and generation [30, 32]. Direct application to REC tasks may still lead to sub-optimal vision-language understanding.
@@
To address these challenges, we propose a novel Multi-Modal Interactive Side-Tuning (M2IST) method that effectively strengthens vision-language alignment and enables parameter- and memory-efficient transfer to REC within the unified interactive side networks (see Fig. 1 (c)). Specifically, we introduce Mixture of Multi-Modal Interactive Side-Adapters (M3ISAs), which incorporate Vision Expert Adapters (VEA), Language Expert Adapters (LEA), and Interaction Expert Adapters (IEA) into the side networks in parallel with the heavy encoders. VEA and LEA transfer pre-trained single-modality knowledge to the REC domain. IEA utilizes a linear layer for weight-sharing between image and text features, enabling progressive interaction between the referring sentence and input image. Such interaction aggregates channel-level vision-language alignment at shallow layers of the model, facilitating deep token-level vision-language alignment in deeper layers for improved performance. This elegant design achieves parameter-, memory-, and time-efficient intra- and inter-modality representation transfer for REC.
@@
We conduct extensive experiments on RefCOCO [33], RefCOCO+ [33], and RefCOCOg [34, 35] to demonstrate the effectiveness and efficiency of M2IST for REC. Experimental results show that M2IST achieves the optimal performance-parameter-memory trade-off compared to most full fine-tuning methods and other PETL methods. By applying our M2IST method, a standard transformer-based REC model can requiring only 2.11% of the tunable parameters, 39.61% of the GPU memory and 63.46% of the fine-tuning time compared to full fine-tuning, while still achieving competitive performance (see Fig. 1). With the sufficient vision-language interaction strengthened by our M3ISAs, our method can accurately locate the referred objects for various complex cases (see Fig. 4).
@@
The rest of this paper is organized as follows. In Section II, we briefly review related work on referring expression comprehension and efficient transfer learning. We then provide a detailed description of our proposed M2IST method and its core component, M3ISA, followed by a discussion of the advantages of M2IST in Section III. In Section IV, we present extensive quantitative and qualitative experiments to analyze the performance and efficiency of M2IST. Finally, we summarize our work in Section V and discuss the limitations and future work in Section VI.
@@ II-A Referring Expression Comprehension
Referring expression comprehension (REC) [1, 2, 3, 7, 4, 9] aims to locate the specific objects in images based on textual descriptions.
Early REC methods [36, 1, 37] follow a two-stage pipeline that first uses a pre-trained object detector (e.g., Faster R-CNN [38]) to generate a set of sparse object proposals, which are then ranked by their similarity to the given textual description. MAttNet [1] uses Faster R-CNN to generate object proposals, then scores them based on the referring expression to find the most relevant object. RvG-Tree [39] builds a relational visual graph to capture object interactions and ranks proposals according to their relevance to the expression, improving grounding accuracy. However, these two-stage REC methods heavily rely on the quality of the object proposals and cannot directly predict the referred object region. One-stage anchor-based methods [2, 40, 41, 42] have been developed to eliminate the proposal generation step, directly predicting object bounding boxes from predefined anchors [43]. FAOA [2] utilizes the YOLOv3 detector [43], integrating it with an encoded language vector to ground the referred regions. MRLN [4] introduces three modules: feature-feature relational learning, feature-task relational learning, and task-task relational learning, to enhance the collaborative learning of REC, effectively reducing prediction inconsistency in multi-task learning. Recently, transformer-based methods [3, 44, 7, 9] have demonstrated superior performance by implicitly modeling cross-modality relationships within a unified architecture. The pioneering work TransVG [3] applies a stack of transformer encoders to perform feature extraction and multi-modal fusion for REC. VGTR [44] employs a transformer encoder-decoder architecture to jointly reason over visual and textual inputs, grounding the referred object without relying on pre-trained detectors or word embeddings. Dynamic M-DETR [9] uses a dynamic multi-modal transformer decoder to adaptively sample visual features and perform text-guided decoding for REC.
As REC models continue to scale up, their performance has shown significant improvements. However, most transformer-based methods, including TransVG [3], VGTR [44], PFOS [7], and Dynamic M-DETR [9], require fully fine-tuning of their pre-trained networks. While these approaches achieve satisfactory performance, they come with a substantial computational cost, demanding larger GPU memory to accommodate the increased number of trainable parameters (see Fig. 1 (a)). To this end, our M2IST aims to provide an efficient fine-tuning strategy for these transformer-based REC models, significantly reducing the computational costs during the fine-tuning stage.
@@ II-B Parameter-efficient Transfer Learning
Recent PETL methods can be classified into two categories. The first category is updating additional parameters in modules inserted into the model [20, 23] or appended to the input data [22, 17, 45]. Adapter [20] incorporates a bottleneck module into each Transformer layer, positioned after both the Multi-Head Attention (MHA) and the Feed-Forward Networks (FFN). AdaptFormer [23] embeds the adapter module parallel to the FFN in each encoder of a Vision Transformer. VPT [22] appends learnable visual vectors to the input sequences (VPT-Shallow) or to the input of each transformer encoder layer (VPT-Deep). The second category involves decomposing weight matrices into two low-rank matrices and updating only the small factorization matrices [21, 46]. As a pioneering work, LoRA [21] integrates a tunable pair of low-rank decomposed weight matrices into each encoder layer of the pre-trained networks. FacT [46] incorporates tunable factorized weight matrices into each layer of the pre-trained networks. There is also growing interest in adapter-based PETL methods for vision-language tasks, including text-image retrieval [18] and text-video retrieval [19]. Most of these methods aim to achieve effective cross-modality interaction while maintaining parameter efficiency.
However, existing PETL methods still face substantial GPU memory consumption during the fine-tuning stage, as gradients must propagate through the heavy pre-trained encoders for REC (see Fig. 1 (b)). Compared with above PETL methods, our M2IST aims to not only reduce the number of updated parameters but also significantly decrease the required GPU memory consumption, thereby providing a more comprehensive efficient fine-tuning strategy.
@@ II-C Memory-efficient Transfer Learning
While existing MEFT methods, particularly side-tuning approaches, have achieved notable results in reducing GPU memory consumption, they are primarily designed for single-modality tasks and fail to effectively bridge pre-trained vision and language models for the challenging task of REC. In this work, inspired by the existing efforts, our M2IST bridges the pre-trained Vision Encoder and Language Encoder through the unified side networks, enabling a parameter-, memory-, and time-efficient transfer to the REC task (see Fig. 1 (c)).
@@ III-A Base Architecture
We apply a standard transformer-based REC model as our base architecture, shown in Fig. 1 (a), which comprises: (1) a Vision Encoder, (2) a Language Encoder, and (3) a Vision-language Encoder.
@@ III-B Multi-Modal Interactive Side-Tuning: M2IST
Given that the pre-trained Vision Encoder and Language Encoder contain rich knowledge and comprise about 95% of the model’s parameters. We first explore two approaches to reduce training overhead:
@@
To address the above issues, we propose Multi-Modal Interactive Side-Tuning (M2IST) that keeps the pre-trained encoders frozen and updates the proposed Mixture of Multi-Modal Interactive Side Adapters (M3ISA) on side networks to facilitate parameter- and memory-efficient fine-tuning for REC, as shown in Fig. 2. Note that we do not show the LayerNorm for simplicity.
@@
The core component of our M2IST is M3ISA (see Fig. 2 (right)), which consists of two distinct adapters, i.e., intra- and inter-modality adapters to effectively and efficiently bridge the pre-trained Vision Encoder and Language Encoder.
The intra-modality adapters include Vision Expert Adapter (VEA) and Language Expert Adapter (LEA), shown as separate blue branch and green branch in Fig. 2 (right). They adhere to a fundamental design [20] for transferring pre-trained single-modality representations to more domain-specific ones. Specifically, both of them consist of a linear down-projection layer $\mathbf{W}_{\text{down}}$, ReLU non-linear activation, and a linear up-projection layer $\mathbf{W}_{\text{up}}$ in sequence. Taking the VEA as a formulaic example, given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, the function of VEA can be formally expressed as:
@@
In the base architecture, the Vision Encoder and Language Encoder are responsible for extracting single-modality features, while only the V-L encoder is tasked with cross-modality fusion through token-level vision-language alignment. However, such the late fusion has been proved insufficient when the referring sentence contains complex semantic information, such as spatial relationships [51]. Therefore, we introduce an inter-modality adapter, named the Interaction Expert Adapter (IEA). Specifically, each IEA shares a set of tokens within the same channel dimensions $C_{i}$ (i.e., Interacted Tokens in Fig. 2 (right)) between the Vision Encoder and Language Encoder, thereby achieving channel-level vision-language alignment during the multi-modal feature extraction stage. As depicted by the entire pink section in Fig. 2 (right), IEA include a unique down-projection layer for vision $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.180,0.459,0.714}C_{v}}\times C_{d}}$ and language $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.439,0.678,0.278}C_{l}}\times C_{d}}$, ReLU activation, an interactive up-projection layer $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times{\color[rgb]{0.929,0.420,0.380}C_{i}}}$, and a unique up-projection layer for vision $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.180,0.459,0.714}C_{v}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$ and language $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.439,0.678,0.278}C_{l}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$, where $C_{v}$, $C_{l}$, and $C_{i}$ represent the vision, language, and interaction channels, respectively. Given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and language tokens $\bm{x}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$, the corresponding down-projection layers first down-sample them to the bottleneck features $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{d}}$ and $\bm{z}_{l}\in\mathbb{R}^{N_{l}\times C_{d}}$. Then, the corresponding up-projection layers and interactive up-projection layer up-sample these bottleneck features and concatenate them within the same modality to obtain the cross-modality features $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$ as:
@@
Algorithm 1 PyTorch-like pseudocode for IEA and M2IST.
# Define the IEA Module
class IEA(nn.Module):
def __init__(self, vis_dim_model, text_dim_model, down_bottle, share_up, inter_up):
self.text_down_proj = nn.Linear(text_dim_model, down_bottle)
self.vis_down_proj = nn.Linear(vis_dim_model, down_bottle)
self.inter_up = nn.Linear(down_bottle, share_up)
self.text_up_proj = nn.Linear(down_bottle, text_dim_model)
self.vis_up_proj = nn.Linear(down_bottle, vis_dim_model)
def forward(self, text_x, vis_x):
vis_down = self.vis_down_proj(vis_x)
text_down = self.text_down_proj(text_x)
vis_inter_up = self.inter_up(vis_down)
text_inter_up = self.inter_up(text_down)
vis_up = self.vis_up_proj(vis_down)
text_up = self.text_up_proj(text_down)
return concat[text_up,text_inter_up], concat[vis_up,vis_inter_up]
IEA_out, LEA_out, VEA_out = [], [], []
# Multi-Modal Interactive Side-Tuning
for i in range(layers):
IEA_out = IEA(text_mha_output,vis_mha_output)
LEA_out = LEA(text_ffn_out)
VEA_out = VEA(vis_ffn_out)
IEA_out.append(IEA_out)
LEA_out.append(LEA_out)
VEA_out.append(VEA_out)
final_text_feature = text_feature + sum(IEA_out) + sum(LEA_out)
final_vis_feature = vis_feature + sum(IEA_out) + sum(VEA_out)
Our IEA enables lightweight adaptation of cross-modality representations without increasing the explicit computational burden of vision-language feature extraction, thereby efficiently bridging the pre-trained vision and language encoders. Our experimental results in Table X further demonstrate that deeper channel-level vision-language alignment yields better performance for REC. Moreover, such alignment will further improve the token-level vision-language alignment in V-L Encoder, leading to better region-sentence understanding in challenging REC scenarios, as analyzed in Section IV-E.
@@
As depicted in Fig. 2 (left), we incorporate a stack of M3ISAs into two side networks that operate in parallel with the pre-trained dual encoders. Specifically, in one encoder layer (both for vision and language), the IEA first receives processed vision/language tokens from the Multi-head Attention (MHA) layers as input and produces adapted, interacted tokens for the vision/language side network. Subsequently, the VEA/LEA take the processed vision/language tokens from the Feed Forward Networks (FFN) as input and generate adapted single-modality tokens for the corresponding side networks. The outputs of the IEA and VEA/LEA are added within the vision/language side networks, along with the original vision/language tokens through skip-connections. After passing through the side networks, the outputs of the vision/language side networks are added to the outputs of the vision/language encoders. During fine-tuning, we keep the pre-trained encoders fixed and update the M3ISAs in the side networks, allowing the pre-trained encoders to act as standalone feature extractors. We present the PyTorch-like pseudocode of the proposed IEA module and M2IST in Algorithm 1 for better understanding.
@@
Following most transformer-based REC methods [3, 9], the training loss function is a combination of the widely used smooth L1 loss and GIoU loss. Specifically, the prediction is donated as $\mathbf{b}=(x,y,w,h)$, and the normalized ground-truth box as $\hat{\mathbf{b}}=(\hat{x},\hat{y},\hat{w},\hat{h})$. The training objective is:
@@
TABLE I: Comparison with efficient tuning methods using the same base architecture. "Param." indicates the number of tunable parameters in the pre-trained encoders. "Mem." denotes the peak GPU memory footprint with batch size 64 during fine-tuning.
We highlight the best and the second results.
(M) (GB) val testA testB val testA testB val-g val-u test-u
PETL Methods:
METL Methods:
DTL [28] 3.22 (2.13%) 15.65 (40.18%) 79.89 80.52 76.33 62.14 66.41 54.49 65.31 66.23 65.94